An airplane covered 20 miles of its route while decreasing its altitude by 47,000 feet. Find the slope of the airplane's line of descent. Round to the nearest hundredth. [Hint: 1 mi = 5280 feet.]
Hint: Should the slope be positive or negative?
Answer: The slope of the airplane's descent is ______________

Answer by vertciel(183)

Hint: The slope should be negative, because the airplane is descending. I assume this is a fighter aircraft because this pilot is popping a lot of passengers' eardrums.
If you draw a quick sketch of this problem, you'll see that the airplane starts at (0, 47000) and descends to (-94,000, 0).
Therefore, slope = rise/run
= (47000 - 0)/(-94000 - 0)
= -47/94 or -0.5